when 12 workers doing 1 work in 12 days then 8 workers doing same work how many days
Found 2 solutions by josgarithmetic, TimothyLamb:
Answer by josgarithmetic(39613)

You expect each worker to work at the same rate as any worker in the group. If 1 worker instead of 12 workers, then he needs twelve times longer time to do the same "work".

Answer by TimothyLamb(4379)

rate:
r = job/(workers*time)
---
rate 12 workers:
1/(12*12)
---
rate 8 workers:
1/(8*t)
---
1/(12*12) = 1/(8*t)
8t = 12*12
t = 12*12/8
t = 18 hours
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php